# HebID: Detecting Social Identities in Hebrew-language Political Text

### v1

@@ 6 Results
To examine which identities tend to go together, we perform a factor analysis on the mean levels of each identity per speaker/respondent (Figure 4). In the three sources, we see a primary dimension dividing identities into two broad groups, a left-wing group (Leftist, Democrat, Honest, Liberal, Palestinian) and a right-wing group (Rightist, Conservative, Zionist, Security-oriented, Capitalist, Ultra-orthodox). Other factors in the Facebook and Knesset data appear to reflect political sub-dimensions of identity pairs that are more tightly coupled, such as Conservative and Rightist, Zionist and Security-oriented, Liberal and Palestinian, and Honest and Democrat. In the survey data, factors beyond the first dimension appear less structured compared to those that arise in the Facebook and Knesset datasets.
Accordingly, examination of the correlations between identities (after aggregating into speakers/respondents) shows that the survey exhibits a mean absolute correlation coefficient of 0.159, weaker than the Facebook sample (0.235) and Knesset data (0.215). For full correlation matrices, see the appendix.
@@
We then examine the gender difference per identity by subtracting the share of each identity among women from that of men, in all data sources. We see that some identities, such as Rightist, Security-oriented, Capitalist, and Ultra-orthodox lean towards men, whereas Socially-oriented significantly lean towards women, on all platforms. While the gender gaps on Facebook and the Knesset are generally in the same direction, several of the survey gender gaps arise in different directions. Thus, Honest in the survey leans significantly towards women, whereas in the Facebook and Knesset data it leans slightly towards men.
@@ 7 Conclusion
In this work, we have introduced HebID, the first publicly available multilabel Hebrew corpus for fine-grained social identity detection in political text, grounded in large-scale survey evidence and expert consultation. Our dataset of 5,536 sentences from Israeli politicians’ Facebook posts is annotated for twelve empirically salient identities, ensuring both linguistic and sociological validity. We benchmarked a range of models—multilabel and single-label encoders as well as 2B–9B-parameter seq2seq LLMs—and demonstrated that Hebrew-tuned seq2seq models (e.g., DictaLM2.0) achieve the highest macro-F1 (0.743), significantly outperforming encoder-only baselines. We further validated cross-genre robustness by applying our best model to 500 Knesset speech excerpts, matching social-media performance (macro-F1 = 0.72).
By linking three complementary sources — Facebook posts, parliamentary speeches, and a national survey, we revealed systematic differences in identity prevalence, temporal surges around elections, bundles of co-occurring identities, and gendered patterns of identity expression. The HebID annotation scheme and corpus, together with our code and models, provides a comprehensive framework for future research on identity discourse in non-English political contexts and paves the way for comparative studies across political systems.
@@
Survey population: The panel survey sampled only Jewish citizens of Israel. Identities and salience patterns may differ among non-Jewish citizens, Palestinians, or immigrant communities.
@@
Cross-genre validation: The Knesset evaluation, while indicative of robustness, is based on a limited sample of 500 human-annotated sentences drawn from Knesset Plenum protocols. A broader evaluation across other legislative bodies or timeframes is needed to fully assess generalisation.
@@ Appendix A Full list of social identities in Survey
Appendix B Identity Definitions: Codebook
@@
This identity focuses on conflict-related positions, not economic or social issues (which are coded separately, e.g., Liberal or Socially-oriented).
@@
Positions regarding state-religion relations are not part of ideological identities (e.g., referring to religious coercion would be considered liberal, not Leftist). Moreover, attacking ultra-Orthodox parties does not constitute a Leftist identity.
@@
This identity focuses on conflict-related positions, not economic or social issues (which are coded separately, e.g., Conservative or Capitalist).
@@
Palestinians and Arab Citizens of Israel. Statements regarding policy issues, worldviews, and ideologies related to Palestinians and Arab-Israelis, including references to Palestinian and Arab-Israeli culture, statements and interviews by Palestinian and Arab-Israeli public leaders related to Palestinians and Arab-Israelis, as well as policy matters concerning these groups (not just security issues).
@@
4.
“No peace without security” statements fall under this identity.
Ultraorthodox (Haredi). References to Jewish ultraorthodox (Haredi) lifestyle, including the Haredian education system, gender segregation in the public sphere, exemption from military service for yeshiva students, charitable activities within the Haredi community, the Haredi religious (Torah) world and tradition preservation, Haredian parties (Shas, United Torah Judaism, Agudat Yisrael), and Haredi rabbinic leadership.
@@
The mere mentioning of Rabies is not enough. Statements must carry a positive tone.
@@
This definition draws on the Declaration of Independence, which defines Israel as the homeland of all Jews while committing to full equality for all its citizens, including minority groups.
@@
This identity includes expressions of dedication to Israel’s well-being or future (“to serve the country,” “for the future of the country”).
3.
This identity does not include general mentions of antisemitism without connection to Israel, references to soldiers unrelated to enlistment or national service.
Appendix C Fine-tuning setup
@@
All experiments utilize AdamW optimizer with a linear scheduler. Default values of hyperparamters are used everywhere except for learning rate (for encoders and for decoder LLMs) and loss type (for encoder models).
Decoder fine-tuning uses QLORA with 4bit quanitzation. LORA settings are rank of 256, and alpha value of 512. LORA layers are attached to all linear levels in the decoder models.
@@
Appendix D Generalization to Parliamentary Speech
@@
Appendix E Data Release
The annotated data is released under cc-by-4.0 license. The data is publicly available on github at https://github.com/guymorlan/pdd/.
Appendix F Annotation
@@
Appendix G Model Sizes
@@ Appendix H Preprocessing
Appendix I Additional Results

### v2

@@ 6 Results
To examine which identities tend to co-occur, we perform a factor analysis on the mean levels of each identity per speaker/respondent (Figure 4). In the three sources, we see a primary dimension dividing identities into two broad groups, a left-wing group (Leftist, Democrat, Honest, Liberal, Palestinian) and a right-wing group (Rightist, Conservative, Zionist, Security-oriented, Capitalist, Ultra-orthodox). Other factors in the Facebook and Knesset data appear to reflect political sub-dimensions of identity pairs that are more tightly coupled, such as Conservative and Rightist, Zionist and Security-oriented, Liberal and Palestinian, and Honest and Democrat. In the survey data, factors beyond the first dimension appear less structured compared to those that arise in the Facebook and Knesset datasets.
Accordingly, examination of the correlations between identities (after aggregating into speakers/respondents) shows that the survey exhibits a mean absolute correlation coefficient of 0.159, weaker than the Facebook sample (0.235) and Knesset data (0.215). For full correlation matrices, see appendix K.
@@
We then examine the gender difference per identity by subtracting the share of each identity among women from that of men, in all data sources. We see that some identities, such as Rightist, Security-oriented, Capitalist, and Ultra-orthodox lean towards men, whereas Socially-oriented significantly leans towards women across all platforms. While the gender gaps on Facebook and the Knesset are generally in the same direction, several of the survey gender gaps arise in different directions. Thus, Honest in the survey leans significantly towards women, whereas in the Facebook and Knesset data it leans slightly towards men.
@@ 7 Conclusion
In this work, we have introduced HebID, the first publicly available multilabel Hebrew corpus for fine-grained social identity detection in political text, grounded in large-scale survey evidence and expert consultation. Our dataset of 5,536 sentences from Israeli politicians’ Facebook posts is annotated for twelve empirically salient identities, ensuring both linguistic and sociological validity. We benchmarked a range of models—multilabel and single-label encoders as well as 2B–9B-parameter decoder LLMs—and demonstrated that Hebrew-tuned decoder models (e.g., DictaLM2.0) achieve the highest macro-F1 (0.743), significantly outperforming encoder-only baselines. We further validated cross-genre robustness by applying our best model to 500 Knesset speech excerpts, matching social-media performance (macro-F1 = 0.72).
By linking three complementary sources — Facebook posts, parliamentary speeches, and a national survey — we revealed systematic differences in identity prevalence, temporal surges around elections, bundles of co-occurring identities, and gendered patterns of identity expression. These findings showcase the potential of our framework to uncover nuanced political dynamics and address substantive questions in the social sciences.
The HebID annotation scheme and corpus provides a comprehensive foundation for future research on identity discourse in non-English political contexts and paves the way for comparative studies across political systems. Specifically, this work provides a new tool for social scientists studying political communication, a new benchmark for NLP researchers working on Hebrew and multilabel classification, and a methodological template for creating similar resources cross-culturally.
@@
Survey population: The panel survey sampled only Jewish citizens of Israel. Identities and salience patterns may differ among non-Jewish citizens, including Palestinians and non-Jewish immigrant communities.
@@
Cross-genre validation: The Knesset evaluation, while indicative of robustness, is based on a limited sample of 500 human-annotated sentences drawn from Knesset Plenum protocols. A broader evaluation across other legislative bodies or timeframes is needed to fully assess generalization.
@@
Acknowledgments
This work was supported by the Israel Science Foundation (Grant no. 2315/18). We thank Vered Porzycki for assistance in conceptualizing the definitions of the annotated identities; Vered Porzycki, Gal Ron, Avishai Green, Dror Markus and Noa Amir for invaluable assistance in a preliminary pilot study; and Effi Levi for his very useful and insightful advice.
@@ Appendix A Full list of social identities in Survey
Appendix B Survey Methodology
The survey we utilize is an online panel survey conducted between January 2019 and April 2021, using a representative sample of the adult Jewish population in Israel. Participants were recruited through iPanel, the largest online survey firm in Israel, and received gift cards as compensation. To ensure representativeness, respondents were selected using demographic quotas based on key population parameters: age, gender, education, geographic location, and religiosity. A total of 1,769 individuals completed the first wave in January 2019, and 878 panelists completed the relevant waves for this study.
To measure identity salience, respondents were asked the following question: "People feel a sense of belonging to some groups in society and a sense of distance from others. Below is a list of groups. Please select up to three groups that best define who you are, groups that make you feel a sense of belonging, identification, and pride, and that you would like to see properly represented in the Knesset."
The possible selection options are the 28 identities appearing in section A of the appendix, a list developed by experts in Israeli politics to ensure relevance and comprehensiveness.
In total, the panel included 14 waves and an additional minor wave in which only several question items were repeated. We utilize 12 of the waves in which the identity salience question appears (11 major waves and the minor wave).
Appendix C Identity Definitions: Codebook
@@
This identity focuses on hawkish-dovish positions, not economic or social issues (which are coded separately, e.g., Liberal or Socially-oriented).
@@
Positions regarding state-religion relations are not considered part of the left/right ideological identities (e.g., referring to religious coercion would be considered liberal, not Leftist).
@@
This identity focuses on hawkish-dovish positions, not economic or social issues (which are coded separately, e.g., Conservative or Capitalist).
@@
Palestinians and Arab Citizens of Israel. Statements regarding policy issues, worldviews, and ideologies that represent or reflect Palestinians and Arab-Israelis, including references to Palestinian and Arab-Israeli culture, statements and interviews by Palestinian and Arab-Israeli public leaders related to Palestinians and Arab-Israelis, as well as policy matters concerning these groups (not just security issues).
@@
Ultra-orthodox (Haredi). References to Jewish ultra-orthodox (Haredi) lifestyle, including the Haredian education system, gender segregation in the public sphere, exemption from military service for yeshiva students, charitable activities within the Haredi community, the Haredi religious (Torah) world and tradition preservation, Haredian parties (Shas, United Torah Judaism, Agudat Yisrael), and Haredi rabbinic leadership.
@@
The mere mentioning of Rabbis is not enough. Statements must carry a positive tone.
@@
This definition draws on the Declaration of Independence, which defines Israel as the homeland of the Jewish people.
@@
This identity includes expressions of dedication to Israel’s future, strength and prosperity (“to serve the country,” “for the future of the country”).
Appendix D Fine-tuning setup
@@
All experiments utilize AdamW optimizer with a linear scheduler. Default values of hyperparameters are used everywhere except for learning rate (for encoders and for decoder LLMs) and loss type (for encoder models).
Decoder fine-tuning uses QLORA with 4bit quantization. LORA settings are rank of 256, and alpha value of 512. LORA layers are attached to all linear levels in the decoder models.
@@
Appendix E Generalization to Parliamentary Speech
@@
Appendix F Data Release
The annotated data is released under cc-by-4.0 license. The data is publicly available on github at https://github.com/guymorlan/hebid/.
Appendix G Annotation
@@
Appendix H Model Sizes
@@ Appendix I Preprocessing
Appendix J External Validity
| CHES variable (2021) | Correlated with | Pearson r | p-value |
| Overall Ideology (Left/Right) | Rightist - Leftist | 0.941 | 0.001 |
| Rightist | 0.880 | 0.002 | |
| Leftist | -0.815 | 0.007 | |
| Social Values (Libertarian/Traditional) | Conservative - Liberal | 0.817 | 0.007 |
| Conservative | 0.431 | 0.247 | |
| Liberal | -0.838 | 0.005 | |
| Civil Liberties vs. Law & Order | Conservative - Liberal | 0.905 | 0.001 |
| Conservative | 0.850 | 0.004 | |
| Liberal | -0.845 | 0.004 | |
| Economic Stance (Left/Right) | Capitalist - Socially-oriented | 0.423 | 0.256 |
| Capitalist | 0.869 | 0.002 | |
| Socially-oriented | 0.012 | 0.976 | |
| State’s Identity (Democratic vs. Jewish) | Ultra-orthodox - Democrat | 0.708 | 0.033 |
| Ultra-orthodox | 0.649 | 0.059 | |
| Democrat | -0.410 | 0.274 | |
| Salience of Reducing Corruption | Honest | 0.664 | 0.051 |
| Stance on Israeli-Palestinian Conflict | Zionist | 0.865 | 0.003 |
| Palestinian | -0.602 | 0.086 | |
| Position on a Palestinian State | Zionist | 0.862 | 0.003 |
| Palestinian | -0.549 | 0.126 | |
| Position on Engagement with Arab World | Security-oriented | 0.710 | 0.032 |
Appendix K Additional Results